In the reaction^7Li(p, d)^6Li, the binding energie
Answer #1

Bindign energy of proton Ep = 7.289 MeV

Q value = E (7Li) + E(p) - [E(6) + E(2D)] = 39.246+7.289 - [31.995+2.225] = 12.315MeV

Eth = -12.315*(1+7)/7 = -14.07 MeV

Eth is greater than Q(in absolute terms) as the entire Kinetic energy is not available to be dissipated as some energy is carried away by the centre of mass.